# Test-Time Iterative Error Correction for Efficient Diffusion Models

### v1

@@ 1 Introduction
Diffusion models ho2020ddpm; rombach2022ldm; nichol2021improved have achieved state-of-the-art generative performance across a wide range of tasks, including image synthesis song2020score; choi2021ilvr; saharia2022palette; tumanyan2023plug; li2022srdiff; saharia2022image; gao2023implicit; avrahami2022blended; kawar2023imagic; meng2021sdedit, text-to-image generation nichol2021glide; ramesh2022hierarchical; saharia2022photorealistic; zhang2023adding, text-to-3D generation lin2023magic3d; luo2021diffusion; poole2022dreamfusion, video generation mei2023vidm, and audio generation huang2022prodiff; zhang2023survey. However, diffusion models typically require hundreds of iterative denoising steps and contain billions of parameters, resulting in high computational costs that hinder deployment in resource-constrained environments. To address this limitation, extensive efficiency techniques have been devoted to developing efficient model techniques for diffusion models song2020ddim; zhang2022gddim; zheng2023fast; liu2025cachequant; shang2023ptq4dm; li2023qdiffusion; li2024qdm.
@@
Among these techniques, network quantization shang2023ptq4dm; li2023qdiffusion; li2024qdm; zheng2024binarydm and feature caching chen2024delta; zou2024accelerating; ma2024deepcache have emerged as two promising approaches. The former reduces data precision to low-bit representations to simultaneously shrink model size and accelerate inference, while the latter caches intermediate features to eliminate redundant computations across diffusion timesteps. Although these methods effectively reduce overhead, they inevitably suffer from approximation errors between the efficient model and the original counterpart, which degrade the generation quality of the model. Prior studies have attempted to mitigate such errors through specialized mechanisms. For example, timestep-wise quantization parameters wang2023towards; liu2024dilatequant have been introduced to capture the dynamics of time-varying activations, while non-uniform caching strategies ma2024deepcache exploit similarity patterns between adjacent timesteps to improve performance. Despite their effectiveness, these methods remain pre-deployment solutions that require the ability to re-execute the model-efficiency pipeline and the original model.
@@
In practice, such requirements often do not hold once a model has been deployed in edge or production environments. On the one hand, reapplying the model-efficiency pipeline and redeploying the model incurs significant engineering overhead, making it impractical in many cases. Also, deployed models are typically immutable due to storage limitations, deployment policies, or system design constraints. On the other hand, after being deployed, the original model is often irretrievable, making re-executing the model-efficiency pipeline unfeasible. For instance, a quantized model may no longer retain its original high-precision weights, making re-quantization infeasible. Inspired by recent advances in test-time scaling snell2024scaling; lightman2023let; muennighoff2025s1; ma2025inference; singhal2025general; prabhudesai2023diffusion; xiao2024beyond, where model behavior is adjusted at inference time without retraining, we ask: Is it possible to improve the performance of an already deployed diffusion model without repeating the model-efficiency pipeline?
@@ 2.1 Diffusion Model
Diffusion models ho2020ddpm; rombach2022ldm consist of forward process and reverse process. In the forward process, given input data distribution $x_{0}\sim q(x)$, diffusion models add a series of Gaussian noise to $x_{0}$ to resulting in a sequence of noisy samples $x_{t},0\leq t\leq T$:
@@
For DDIM song2020ddim, the $\hat{\beta}_{t}=0$ and $\hat{\mu}_{\theta,t}$ is defined by:
@@ 2.2 Efficient Diffusion
Significant efforts have been devoted to developing efficient diffusion models, which can be broadly categorized into two types: temporal efficiency and structural efficiency liu2025cachequant. Temporal efficiency methods focus on reducing the number of sampling timesteps. For instance, methods such as DDIM song2020ddim and GDDIM zhang2022gddim modify the denoising equations to enable fewer sampling steps, while others zheng2023fast; shih2023parallel accelerate inference by performing multiple denoising timesteps in parallel. DistriFusion li2024distrifusion further improves efficiency by dividing input patches across multiple GPUs. Additionally, fast solvers for diffusion liu2022pseudo; dockhorn2022genie; yin2024one; yin2024improved; lu2022dpm; lu2022dpmplusplus have been proposed to reduce the computational overhead associated with sampling. Structural efficiency methods, on the other hand, aim to reduce model complexity and are complementary to temporal efficiency techniques, making them increasingly popular. For example, network quantization has emerged as a key approach for efficient diffusion model li2024qdm; zheng2024binarydm; yang2024analysis; deng2025vq4dit; zhao2024viditq; Huang_2024_CVPR; chen2024q; dong2025ditas. Quantization-aware training (QAT)li2024qdm; zheng2024binarydm restores model performance at low bit-widths but incurs significant training overhead. In contrast, post-training quantization (PTQ)shang2023ptq4dm; li2023qdiffusion; wu2024ptq4dit requires minimal computational resources and small calibration datasets, making it more suitable for scenarios with limited resources. PTQ4DM shang2023ptq4dm and Q-Diffusion li2023qdiffusion introduced time-aware calibration to realize quantized diffusion models. Subsequent works have refined these methods by addressing quantization error, feature re-balancing, calibration data collection, model reconstruction, mixed-precision strategies, and so on he2024ptqd; huang2024tfmq; wang2024apqdm; wang2024quest; feng2025mpq; liu2024enhanced; sun2024tmpq; wuptq4dit. Moreover, techniques such as LoRA-based optimization hu2021lora have been applied to further enhance quantized diffusion models lisvdquant; he2023efficientdm. In addition to quantization, feature caching has also gained attention chen2024delta; zou2024accelerating; ma2024deepcache; wimbauer2024cache; liu2025reusing; liu2024timestep; ma2024learningtocache. These methods aim to reduce inference time by reusing pre-computed features across diffusion steps. Recent studies, such as CacheQuant liu2025cachequant, combine caching with quantization to achieve even greater efficiency. Other structural efficiency techniques include token pruning fang2024tinyfusion and sparsity fan2025sq. In this work, we focus on mitigating the errors introduced by structural efficiency methods.
@@ 3.2 Iterative Error Correction
According to Banach’s fixed-point theorem Banach1922, the iterative procedure converges to a unique fixed point $x_{t-1}^{*}$ if the mapping $G(x)$ is a contraction mapping, i.e., there exists a Lipschitz constant $0<L<1$ such that:
@@
To guarantee convergence, it is necessary to satisfy $0<L<1$. Since $B_{t}<0$, an appropriately chosen positive $\lambda$ reduces the Lipschitz constant $L$ via the term $\lambda B_{t}J_{t}$, ensuring $L<1$. Empirically, as shown in Fig. 1(b), we observe that setting $\lambda$ within the range [0.1, 0.7] consistently ensures that $\|\nabla G(x)\|<1$ across all timesteps. In this paper, we set $\lambda$ to 0.5 as a practical choice based on these observations. Moreover, since $G(x)$ is continuously differentiable, Eq. 14 holds by the Mean Value Inequality for vector-valued functions munkres2018analysis, confirming that $G$ is a contraction mapping and IEC can converge to the fixed-point solution.
@@ 4.1 Experiment Settings
All experiments are conducted using PyTorch on an NVIDIA 3090 GPU. To evaluate the effectiveness of IEC, we apply it to various diffusion models, including DDPM, LDM, and Stable Diffusion song2020ddim; rombach2022ldm, as well as efficiency techniques such as timestep-wise network quantization liu2024dilatequant, Deepcache ma2024deepcache, and CacheQuant liu2025cachequant, a hybrid technique combining quantization and feature caching. For model quantization, we adopt channel-wise quantization for weights and layer-wise quantization for activations, considering both W4A8 and W8A8 cases. For W4A8, local reconstruction is performed to enhance performance. For feature caching, following CacheQuant ma2024deepcache, the cached blocks are set to last 3, 1, and 1 blocks for DDIM, LDM, and Stable Diffusion, respectively. We evaluate IEC on several widely used datasets, including CIFAR-10 krizhevsky2009learning, LSUN-Churchs yu2015lsun, LSUN-Bedrooms yu2015lsun, ImageNet deng2009imagenet, and MS-COCO lin2014microsoft. For Stable Diffusion, we generate 5,000 images using MS-COCO captions as prompts, following the protocols in ma2024deepcache; chen2024delta; wu2024ptq4dit. For all other datasets, 50,000 images are generated to evaluate the quality of image synthesis. The evaluation metrics include FID heusel2017fid, Inception Score (IS) salimans2016IS, and CLIP Score (evaluated on ViT-g/14) hessel2021clipscore. For quantization-based methods, IEC is applied at every timestep. For DeepCache and CacheQuant, IEC is applied only to the non-cached timesteps, except for MS-COCO, where IEC is only applied at the first timestep. Note that IEC is applied solely to the efficient diffusion models and does not interfere with the baseline model efficiency pipelines.
Table 1: Results of combining IEC with timestep-wise quantization liu2024dilatequant on CIFAR-10, LSUN-Churchs, and LSUN-Bedrooms datasets.
CIFAR-10 32 $\times$ 32
(T=100)
LSUN-Churchs 256 $\times$ 256
(T=100)
@@
LSUN-Bedrooms 256 $\times$ 256
(T=100)
@@
Table 2: Results of combining IEC with Deepcache ma2024deepcache on CIFAR-10, LSUN-Churchs, and LSUN-Bedrooms datasets.
CIFAR 32 $\times$ 32
(T=100)
LSUN-Churchs 256 $\times$ 256
(T=100)
LSUN-Bedrooms 256 × 256
(T=100)
@@ 4.2 Ablation Study
Table 3: Results of combining IEC with CacheQuant liu2025cachequant (Denoted as CacheQ) on CIFAR-10, LSUN-Churchs, and LSUN-Bedrooms datasets.
W/A
CIFAR-10
32 $\times$ 32 (T=100)
LSUN-Churchs
256 $\times$ 256 (T=100)
@@ 32 $\times$ 32 (T=100)
W/A
LSUN-Bedrooms
256 $\times$ 256 (T=100)
@@ 4.3 Main Results
Evaluation on IEC on Feature Caching. The performance of IEC on DeepCache ma2024deepcache is shown in Tab. 2, demonstrating consistent improvements. For example, on CIFAR-10, IEC reduces FID from 4.70 to 3.96 when N = 3. Similar trends are observed for larger N, with FID reduced from 17.21 to 14.58 when N = 15. On LSUN-Churchs, IEC reduces FID from 25.27 to 22.42 at N = 15, and on LSUN-Bedrooms, from 26.09 to 16.91 at N = 10.
Combined With Quantization-Caching. The results of IEC on CacheQuant liu2025cachequant, which integrates quantization and feature caching, are shown in Tab. 3 and Tab. 4. On CIFAR-10, IEC achieves consistent improvements. For instance, under W8A8 with N = 3, FID is reduced from 4.61 to 3.93. With N = 15, FID improves from 13.42 to 10.77. Similar improvements are observed under W4A8. For example, at N = 10, FID improves from 11.36 to 10.69. On LSUN-Churchs, for W8A8 and N = 3, IEC reduces FID from 3.66 to 3.39. At N = 15, FID drops from 9.47 to 6.90. Under W4A8, IEC improves FID by 0.23, 0.45, 1.06, and 1.78 for N = 3, 5, 10, and 15, respectively. On LSUN-Bedrooms, IEC also provides consistent gains. For example, under W8A8 with N = 2, FID improves from 8.85 to 7.51. As shown in Tab. 4, on ImageNet, IEC enhances performance across all baselines. For instance, under W8A8 with N = 10, FID improves from 4.68 to 4.15 and IS from 184.38 to 196.20. At N = 20, FID decreases from 7.21 to 6.53, and IS increases from 160.68 to 169.71. Under W4A8, IEC also improves performance. For example, at N = 10, FID improves from 6.90 to 6.50 and IS from 158.27 to 161.86, demonstrating IEC’s robustness in low-precision scenarios. On MS-COCO, IEC also brings improvements. For example, under W8A8 with N = 10, FID improves from 23.65 to 23.36, IS from 36.71 to 37.02, and CLIP Score from 26.41 to 26.45. At N = 5, FID improves from 23.74 to 22.83, IS from 39.81 to 40.91, and CLIP Score from 26.87 to 26.94.
@@
Table 4: Results of combining IEC with CacheQuant liu2025cachequant (Denoted as CacheQ) on ImageNet and MS-COCO datasets.
@@ A.1 Overhead Discussion
Table 5 presents the overhead and FID when combining our IEC with timestep-wise quantization liu2024dilatequant and Deepcache ma2024deepcache on the CIFAR-10 dataset. Here, we measure the time overhead by comparing it with the baseline without IEC. For W8A8 quantization, IEC achieves substantial improvements in FID with minimal overhead. For instance, selectively applying IEC to $\pm$1/10 or $\pm$1/20 of the steps achieves improved FID of 3.88 and 3.97 with only 20% and 10% overhead, respectively. In the case of DeepCache, the IEC is only applied to the non-cached timesteps, thereby its overhead is minimal. For example, applying IEC to all timesteps yields a FID of 7.77 with only 14% overhead, while applying IEC to $\pm$1/10 or $\pm$1/20 reduces the FID to 9.58 and 9.55, respectively, with overheads as low as 2.8% and 1.4%. These results demonstrate that IEC provides a flexible trade-off between efficiency and generation quality, making it easy to use and suitable for real-world applications.
Table 5: Overhead of combining IEC with timestep-wise quantization liu2024dilatequant and Deepcache (N=10) ma2024deepcache on CIFAR-10.

### v2

@@ 1 Introduction
Diffusion models Ho et al. (2020); Rombach et al. (2022); Nichol and Dhariwal (2021) have achieved state-of-the-art generative performance across a wide range of tasks, including image synthesis Song et al. (2020b); Choi et al. (2021); Saharia et al. (2022a); Tumanyan et al. (2023); Li et al. (2022); Saharia et al. (2022c); Gao et al. (2023); Avrahami et al. (2022); Kawar et al. (2023); Meng et al. (2021), text-to-image generation Nichol et al. (2021); Ramesh et al. (2022); Saharia et al. (2022b); Zhang et al. (2023b), text-to-3D generation Lin et al. (2023); Luo and Hu (2021); Poole et al. (2022), video generation Mei and Patel (2023), and audio generation Huang et al. (2022); Zhang et al. (2023a). However, diffusion models typically require hundreds of iterative denoising steps and contain billions of parameters, resulting in high computational costs that hinder deployment in resource-constrained environments. To address this limitation, extensive efficiency techniques have been devoted to developing efficient model techniques for diffusion models Song et al. (2020a); Zhang et al. (2022); Zheng et al. (2023); Liu et al. (2025b); Shang et al. (2023); Li et al. (2023; 2024c).
@@
Among these techniques, network quantization Shang et al. (2023); Li et al. (2023; 2024c); Zheng et al. (2024) and feature caching Chen et al. (2024b); Zou et al. (2024); Ma et al. (2024b) have emerged as two promising approaches. The former reduces data precision to low-bit representations to simultaneously shrink model size and accelerate inference, while the latter caches intermediate features to eliminate redundant computations across diffusion timesteps. Although these methods effectively reduce overhead, they inevitably suffer from approximation errors between the efficient model and the original counterpart, which degrade the generation quality of the model. Prior studies have attempted to mitigate such errors through specialized mechanisms. For example, timestep-wise quantization parameters Wang et al. (2023); Liu et al. (2024b) have been introduced to capture the dynamics of time-varying activations, while non-uniform caching strategies Ma et al. (2024b) exploit similarity patterns between adjacent timesteps to improve performance. Despite their effectiveness, these methods remain pre-deployment solutions that require the ability to re-execute the model-efficiency pipeline and the original model.
@@
In practice, such requirements often do not hold once a model has been deployed in edge or production environments. On the one hand, reapplying the model-efficiency pipeline and redeploying the model incurs significant engineering overhead, making it impractical in many cases. Also, deployed models are typically immutable due to storage limitations, deployment policies, or system design constraints. On the other hand, after being deployed, the original model is often irretrievable, making re-executing the model-efficiency pipeline unfeasible. For instance, a quantized model may no longer retain its original high-precision weights, making re-quantization infeasible. Inspired by recent advances in test-time scaling Snell et al. (2024); Lightman et al. (2023); Muennighoff et al. (2025); Ma et al. (2025); Singhal et al. (2025); Prabhudesai et al. (2023); Xiao and Snoek (2024), where model behavior is adjusted at inference time without retraining, we ask: Is it possible to improve the performance of an already deployed diffusion model without repeating the model-efficiency pipeline?
@@ 2.1 Diffusion Model
Diffusion models Ho et al. (2020); Rombach et al. (2022) consist of forward process and reverse process. In the forward process, given input data distribution $x_{0}\sim q(x)$, diffusion models add a series of Gaussian noise to $x_{0}$ to resulting in a sequence of noisy samples $x_{t},0\leq t\leq T$:
@@
For DDIM Song et al. (2020a), the $\hat{\beta}_{t}=0$ and $\hat{\mu}_{\theta,t}$ is defined by:
@@ 2.2 Efficient Diffusion
Significant efforts have been devoted to developing efficient diffusion models, which can be broadly categorized into two types: temporal efficiency and structural efficiency Liu et al. (2025b). Temporal efficiency methods focus on reducing the number of sampling timesteps. For instance, methods such as DDIM Song et al. (2020a) and GDDIM Zhang et al. (2022) modify the denoising equations to enable fewer sampling steps, while others Zheng et al. (2023); Shih et al. (2023) accelerate inference by performing multiple denoising timesteps in parallel. DistriFusion Li et al. (2024a) further improves efficiency by dividing input patches across multiple GPUs. Additionally, fast solvers for diffusion Liu et al. (2022); Dockhorn et al. (2022); Yin et al. (2024b; a); Lu et al. (2022a; b) have been proposed to reduce the computational overhead associated with sampling. Structural efficiency methods, on the other hand, aim to reduce model complexity and are complementary to temporal efficiency techniques, making them increasingly popular. For example, network quantization has emerged as a key approach for efficient diffusion model Li et al. (2024c); Zheng et al. (2024); Yang et al. (2024); Deng et al. (2025); Zhao et al. (2024); Huang et al. (2024a); Chen et al. (2024a); Dong and Zhang (2025). Quantization-aware training (QAT)Li et al. (2024c); Zheng et al. (2024) restores model performance at low bit-widths but incurs significant training overhead. In contrast, post-training quantization (PTQ)Shang et al. (2023); Li et al. (2023); Wu et al. (2024a) requires minimal computational resources and small calibration datasets, making it more suitable for scenarios with limited resources. PTQ4DM Shang et al. (2023) and Q-Diffusion Li et al. (2023) introduced time-aware calibration to realize quantized diffusion models. Subsequent works have refined these methods by addressing quantization error, feature re-balancing, calibration data collection, model reconstruction, mixed-precision strategies, and so on He et al. (2024); Huang et al. (2024b); Wang et al. (2024a; b); Feng et al. (2025); Liu et al. (2024c); Sun et al. (2024); Wu et al. (2024b). Moreover, techniques such as LoRA-based optimization Hu et al. (2021) have been applied to further enhance quantized diffusion models Li et al. (2024b); He et al. (2023). In addition to quantization, feature caching has also gained attention Chen et al. (2024b); Zou et al. (2024); Ma et al. (2024b); Wimbauer et al. (2024); Liu et al. (2025a; 2024a); Ma et al. (2024a). These methods aim to reduce inference time by reusing pre-computed features across diffusion steps. Recent studies, such as CacheQuant Liu et al. (2025b), combine caching with quantization to achieve even greater efficiency. Other structural efficiency techniques include token pruning Fang et al. (2024) and sparsity Fan et al. (2025). In this work, we focus on mitigating the errors introduced by structural efficiency methods.
@@ 3.2 Iterative Error Correction
According to Banach’s fixed-point theorem Banach (1922), the iterative procedure converges to a unique fixed point $x_{t-1}^{*}$ if the mapping $G(x)$ is a contraction mapping, i.e., there exists a Lipschitz constant $0<L<1$ such that:
@@
To guarantee convergence, it is necessary to satisfy $0<L<1$. Since $B_{t}<0$, an appropriately chosen positive $\lambda$ reduces the Lipschitz constant $L$ via the term $\lambda B_{t}J_{t}$, ensuring $L<1$. Empirically, as shown in Fig. 1(b), we observe that setting $\lambda$ within the range [0.1, 0.7] consistently ensures that $\|\nabla G(x)\|<1$ across all timesteps. In this paper, we set $\lambda$ to 0.5 as a practical choice based on these observations. Moreover, since $G(x)$ is continuously differentiable, Eq. 14 holds by the Mean Value Inequality for vector-valued functions Munkres (2018), confirming that $G$ is a contraction mapping and IEC can converge to the fixed-point solution.
@@ 4.1 Experiment Settings
All experiments are conducted using PyTorch on an NVIDIA 3090 GPU. To evaluate the effectiveness of IEC, we apply it to various diffusion models, including DDPM, LDM, and Stable Diffusion Song et al. (2020a); Rombach et al. (2022), as well as efficiency techniques such as timestep-wise network quantization Liu et al. (2024b), Deepcache Ma et al. (2024b), and CacheQuant Liu et al. (2025b), a hybrid technique combining quantization and feature caching. For model quantization, we adopt channel-wise quantization for weights and layer-wise quantization for activations, considering both W4A8 and W8A8 cases. For W4A8, local reconstruction is performed to enhance performance. For feature caching, following CacheQuant Ma et al. (2024b), the cached blocks are set to last 3, 1, and 1 blocks for DDIM, LDM, and Stable Diffusion, respectively. We evaluate IEC on several widely used datasets, including CIFAR-10 Krizhevsky et al. (2009), LSUN-Churchs Yu et al. (2015), LSUN-Bedrooms Yu et al. (2015), ImageNet Deng et al. (2009), and MS-COCO Lin et al. (2014). For Stable Diffusion, we generate 5,000 images using MS-COCO captions as prompts, following the protocols in Ma et al. (2024b); Chen et al. (2024b); Wu et al. (2024a). For all other datasets, 50,000 images are generated to evaluate the quality of image synthesis. The evaluation metrics include FID Heusel et al. (2017), Inception Score (IS) Salimans et al. (2016), and CLIP Score (evaluated on ViT-g/14) Hessel et al. (2021). For quantization-based methods, IEC is applied at every timestep. For DeepCache and CacheQuant, IEC is applied only to the non-cached timesteps, except for MS-COCO, where IEC is only applied at the first timestep. Note that IEC is applied solely to the efficient diffusion models and does not interfere with the baseline model efficiency pipelines.
Table 1: Results of combining IEC with timestep-wise quantization Liu et al. (2024b) on CIFAR-10, LSUN-Churchs, and LSUN-Bedrooms datasets.
CIFAR-10 32 $\times$ 32 (T=100) LSUN-Churchs 256 $\times$ 256 (T=100)
@@
LSUN-Bedrooms 256 $\times$ 256 (T=100)
@@
Table 2: Results of combining IEC with Deepcache Ma et al. (2024b) on CIFAR-10, LSUN-Churchs, and LSUN-Bedrooms datasets.
CIFAR 32 $\times$ 32 (T=100) LSUN-Churchs 256 $\times$ 256 (T=100) LSUN-Bedrooms 256 × 256 (T=100)
@@ 4.2 Ablation Study
Table 3: Results of combining IEC with CacheQuant Liu et al. (2025b) (Denoted as CacheQ) on CIFAR-10, LSUN-Churchs, and LSUN-Bedrooms datasets.
W/A CIFAR-10 32 $\times$ 32 (T=100) LSUN-Churchs 256 $\times$ 256 (T=100)
@@
W/A LSUN-Bedrooms 256 $\times$ 256 (T=100)
@@ 4.3 Main Results
Evaluation on IEC on Feature Caching. The performance of IEC on DeepCache Ma et al. (2024b) is shown in Tab. 2, demonstrating consistent improvements. For example, on CIFAR-10, IEC reduces FID from 4.70 to 3.96 when N = 3. Similar trends are observed for larger N, with FID reduced from 17.21 to 14.58 when N = 15. On LSUN-Churchs, IEC reduces FID from 25.27 to 22.42 at N = 15, and on LSUN-Bedrooms, from 26.09 to 16.91 at N = 10.
Combined With Quantization-Caching. The results of IEC on CacheQuant Liu et al. (2025b), which integrates quantization and feature caching, are shown in Tab. 3 and Tab. 4. On CIFAR-10, IEC achieves consistent improvements. For instance, under W8A8 with N = 3, FID is reduced from 4.61 to 3.93. With N = 15, FID improves from 13.42 to 10.77. Similar improvements are observed under W4A8. For example, at N = 10, FID improves from 11.36 to 10.69. On LSUN-Churchs, for W8A8 and N = 3, IEC reduces FID from 3.66 to 3.39. At N = 15, FID drops from 9.47 to 6.90. Under W4A8, IEC improves FID by 0.23, 0.45, 1.06, and 1.78 for N = 3, 5, 10, and 15, respectively. On LSUN-Bedrooms, IEC also provides consistent gains. For example, under W8A8 with N = 2, FID improves from 8.85 to 7.51. As shown in Tab. 4, on ImageNet, IEC enhances performance across all baselines. For instance, under W8A8 with N = 10, FID improves from 4.68 to 4.15 and IS from 184.38 to 196.20. At N = 20, FID decreases from 7.21 to 6.53, and IS increases from 160.68 to 169.71. Under W4A8, IEC also improves performance. For example, at N = 10, FID improves from 6.90 to 6.50 and IS from 158.27 to 161.86, demonstrating IEC’s robustness in low-precision scenarios. On MS-COCO, IEC also brings improvements. For example, under W8A8 with N = 10, FID improves from 23.65 to 23.36, IS from 36.71 to 37.02, and CLIP Score from 26.41 to 26.45. At N = 5, FID improves from 23.74 to 22.83, IS from 39.81 to 40.91, and CLIP Score from 26.87 to 26.94.
@@
Table 4: Results of combining IEC with CacheQuant Liu et al. (2025b) (Denoted as CacheQ) on ImageNet and MS-COCO datasets.
@@ A.1 Overhead Discussion
Table 5 presents the overhead and FID when combining our IEC with timestep-wise quantization Liu et al. (2024b) and Deepcache Ma et al. (2024b) on the CIFAR-10 dataset. Here, we measure the time overhead by comparing it with the baseline without IEC. For W8A8 quantization, IEC achieves substantial improvements in FID with minimal overhead. For instance, selectively applying IEC to $\pm$1/10 or $\pm$1/20 of the steps achieves improved FID of 3.88 and 3.97 with only 20% and 10% overhead, respectively. In the case of DeepCache, the IEC is only applied to the non-cached timesteps, thereby its overhead is minimal. For example, applying IEC to all timesteps yields a FID of 7.77 with only 14% overhead, while applying IEC to $\pm$1/10 or $\pm$1/20 reduces the FID to 9.58 and 9.55, respectively, with overheads as low as 2.8% and 1.4%. These results demonstrate that IEC provides a flexible trade-off between efficiency and generation quality, making it easy to use and suitable for real-world applications.
Table 5: Overhead of combining IEC with timestep-wise quantization Liu et al. (2024b) and Deepcache (N=10) Ma et al. (2024b) on CIFAR-10.